OpenStudy (anonymous):

Write an equation of a line in standard form with slope of -8 passing through the point (-10,2)

OpenStudy (anonymous):

@kierg10, so this one is standard form if you want to take it away :)

OpenStudy (anonymous):

lol :) my fave! y=mx+b m=-8 y=2 x=-10 2=-8(-10)+b b=-78 y=-8x-78

OpenStudy (anonymous):

:D

OpenStudy (anonymous):

@cacique, see how he got the answer?

OpenStudy (anonymous):

You plug in the x, y, and m into the equation and solve for b

OpenStudy (anonymous):

Satndard form is \(Ax + By = C\), that's the slope intercept form, \(y=mx+b\)

OpenStudy (anonymous):

*standard

OpenStudy (anonymous):

Oh that is good to know. Point-slope form is y-y1=m(x-x1)

OpenStudy (anonymous):

I should know standard form. I use it all of the time in my math problems

OpenStudy (anonymous):

heh true XD

OpenStudy (anonymous):

@Data_LG2, I would give you a medal if I could :)

OpenStudy (anonymous):

no worries ;) just change the answer into standard form

OpenStudy (anonymous):

thank you